OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. 2x2 - 10x + 7 = 0

OpenStudy (anonymous):

\[\frac{ -10\sqrt{-10^{2}-4(2)(7)} }{ 2(2) }\]

OpenStudy (anonymous):

is this right?

OpenStudy (abdullahm):

Almost...

OpenStudy (abdullahm):

The quadratic formula is \(\sf\LARGE x=\frac{ -b\pm \sqrt{b^2-4ac}}{2a}\) where the equation is: \(\sf\Large ax^2+bx+c=0\)

OpenStudy (abdullahm):

Remember, your b is negative and we are doing -B in the first part. So you should get \(\sf -(-10)\)

OpenStudy (anonymous):

\[\frac{ 10\pm \sqrt{-10^{2}-4(2)(7)} }{ 2(2) }\] so like that maybe

OpenStudy (abdullahm):

correct

OpenStudy (abdullahm):

but just remember it's \(\sf (-10)^2\) not \(\sf -(10^2)\)

OpenStudy (anonymous):

ok thanks :)